Find domaion
34. x- $(x – 3)-3 - 2-$(x – 3)-$ (22 – 3x + 2) > 0
Answer #1

x^{-1/3}(x-3)^{-2/3} -x^{-4/3}(x-3)^{-5/3}(x^2-3x+2) \geq 0

taking common x^{-4/3} and (x-3)^{-5/3} from expression

x^{-4/3}(x-3)^{-5/3} (x(x-3)-(x^2-3x+2)) \geq 0

x^{-4/3}(x-3)^{-5/3} (x^2-3x-x^2+3x-2) \geq 0

x^{-4/3}(x-3)^{-5/3} (-2) \geq 0

\implies x^{-4/3}(x-3)^{-5/3} \leq 0

we know   x^{-4/3} \geq 0 since there is even power

\implies (x-3)^{-5/3} \leq 0

Domain is x \in (-\infty, 3]